### v2

# Large-Scale Continual Scheduling and Execution for Dynamic Distributed Satellite Constellation Observation Allocation

Abstract.
The size and capabilities of Earth-observing satellite constellations are rapidly increasing. Leveraging distributed onboard control, we can enable novel time-sensitive measurements and responses. However, deploying autonomy to large multiagent satellite systems necessitates algorithms with efficient computation and communication. We tackle this challenge and propose new, online algorithms for large-scale dynamic distributed constraint optimization problems (DDCOP). We present the Dynamic Multi-Satellite Constellation Observation Scheduling Problem (DCOSP), a new formulation of DDCOPs that models integrated scheduling and execution. We construct an omniscient offline algorithm to compute the novel optimality condition of DCOSP and present the Dynamic Incremental Neighborhood Stochastic Search (D-NSS) algorithm, an incomplete online decomposition-based DDCOP approach. We show through simulation that D-NSS converges to near-optimal solutions and outperforms DDCOP baselines in terms of solution quality, computation time, and message volume. Our work forms the foundation of the largest in-space demonstration of distributed multiagent AI to date: the NASA FAME mission.
Key words and phrases:
Distributed constraint optimization, scheduling, satellite operations
1. Introduction
There has been a proliferation of Earth-observing spacecraft in recent years, including advancements in their capabilities to act as autonomous agents. Reduced launch costs have led to constellations composed of hundreds or thousands of spacecraft that can monitor Earth phenomena newspace. Large observation systems result in shorter revisit times to the target observation locations. Reduced revisit times are crucial for rapid responses to dynamic events such as natural disasters. New spacecraft also possess edge hardware capable of performing more intensive computation onboard, including neural network execution and even planning eo1; nos2024; zilberstein-isairas-2024; dt-spaceops-2025; rabideau2025planning. The advancement of inter-satellite links (ISL) has enabled persistent communications between spacecraft and stations on the ground.
These capabilities support observation campaigns that require time-sensitive and coordinated measurements. An example is global monitoring of all volcanic activity or flooding events. However, without consistent observation and the ability to react to dynamic events, key measurements of these processes may be missed. For example, wildfire monitoring requires 30-minute updates to be useful to ground responders wildfireworkshop. Centralized scheduling on the ground suffers from latency and may not be able to meet these timing constraints. Therefore, to successfully tackle these campaigns, we require large-scale observation systems with reduced revisit times and distributed autonomy so that agents can operate online without ground control.
This work focuses on distributed online scheduling that can efficiently coordinate the actions of hundreds or thousands of spacecraft continually as problem dynamics change. Coordinating the actions of a large-scale constellation requires reasoning about agents with varying capabilities, constraints, and visibility of Earth targets while managing limited computational resources. Satellites share CPU and RAM with other critical flight software. Large volumes of communication in space are also unreliable and may even carry a financial cost leo-comms. These limitations make solving a static problem challenging. However, any operational solution must solve dynamic problems that change over time, which further increases the complexity. In these scenarios, the satellite constellation is fixed, yet the observation requests change. These changing observation requests alter the constraints of the problem. We desire continual scheduling solutions that are lightweight yet can effectively handle these problem dynamics.
Operational satellite observation scheduling uses centralized paradigms shah2019scheduling, and many research efforts have concentrated on centralized solutions boerkoel2021efficientsatellite; nag2018scheduling. Centralized approaches may be insufficient for dynamic situations that require real-time response due to latencies to the ground. Centralized approaches are also vulnerable to single-point failures that would result in a non-operational constellation.
We model a large-scale satellite constellation as a multi-agent system (MAS) and focus on continual decentralized scheduling that optimizes observation completion during an overlapping scheduling and execution horizon. We present the Dynamic Multi-Satellite Constellation Observation Scheduling Problem (DCOSP), which is a dynamic distributed constraint optimization problem (DDCOP). DCOSP is a novel application of a DDCOP for dynamic satellite observation scheduling, and extends DDCOPs in multiple ways. The assumptions of DCOSP differ from those of a DDCOP to reflect real-world constraints. We assume agents are aware of the existence of other agents, but do not have access to detailed capabilities or state information. This assumption allows our approach to apply to scenarios with intermittent connectivity, limited bandwidth, or security-mandated communication restrictions. We also formulate a novel optimality condition for a DDCOP that takes into account integrated scheduling and execution. In addition, DDCOP solutions tend to rely heavily on computation and communication, especially when problem dynamics are volatile, which prevents application to DCOSP. These challenges make current DDCOP solvers insufficient for DCOSP.
DCOSP problem instances are much larger than typical DDCOP problems examined in previous literature. DCOSP instances consist of millions of variables that change over time. Agents must react to problem dynamics while expending limited time, memory, and communication. We extend the Neighborhood Stochastic Search (NSS) algorithm to a dynamic variation referred to as Dynamic Incremental Neighborhood Stochastic Search (D-NSS). NSS decomposes the global problem into smaller sub-problems, producing global solutions more efficiently. D-NSS uses the same foundation and leverages repairing previous solutions to efficiently handle problem dynamics.
The contributions of this work are
  1. (1)
    formulating the real-world application of dynamic satellite scheduling as a DDCOP with a unique optimality condition that models task execution,
  2. (2)
    constructing an omniscient, offline optimal solution to the dynamic satellite scheduling problem, and
  3. (3)
    presenting the Dynamic Incremental Neighborhood Stochastic Search algorithm, a scalable incomplete DDCOP approach.
We evaluate the effectiveness of the approaches on large-scale real-world scenarios as well as analyze the challenges of deploying existing DDCOP solutions to this problem. DCOSP and D-NSS will be leveraged in the largest in-space demonstration of multi-agent AI to date, beginning in 2026. The NASA FAME demonstration involves over 60 participating spacecraft that will dynamically coordinate to observe Earth phenomena fame-spaceops-2025.
2. Related Work
2.1. Satellite Observation Scheduling
Satellite observation scheduling is typically framed as an optimization problem that involves geometric reasoning, downlink scheduling, and constraint-based task allocation. The majority of research efforts and operational work have focused on centralized solutions to satellite observation scheduling globus; augenstein2016optimal; nag2018scheduling; he2018improved; shah2019scheduling; aeossp; squillaci2021managing; boerkoel2021efficientsatellite; eddy2021maximum; squillaci2023scheduling; sat-task-plan-large-areas. There is limited work on decentralized scheduling approaches, and these mainly focus on static problems. Examples include auction-based methods picardauction; phillipsauction and heuristic search-based methods shreya; bonnet2007collaboration; bonnet2008coordination; zilberstein2025decentralized. The work of Zilberstein, Rao, Salis, and Chien proposed the Multi-Satellite Constellation Observation Scheduling Problem (COSP), a DCOP formulation of observation allocation zilberstein2025decentralized. We extend this work by formulating the Dynamic Multi-Satellite Constellation Observation Scheduling Problem (DCOSP), which is a novel DDCOP formulation of the problem.
2.2. Dynamic Distributed Constraint Optimization
Distributed constraint optimization problems (DCOP) have modeled applications including mobile sensor teams ca-dcop-mst, smart grids smartgrid, and satellite scheduling zilberstein2025decentralized. Solutions to distributed constraint optimization problems tend to be intensive in computation and communication, making deployment to agents with limited computation challenging. Optimal solutions have exponential complexities syncbb; adopt; gershman2009asynchronous; dpop; mailler2004solving. Incomplete DCOP algorithms are more efficient, yet typically rely on agents communicating with all neighboring agents in the constraint graph, resulting in large complexities when constraint graphs are fully connected mgm; dsa; maxsum; dgibs. The Neighborhood Stochastic Search (NSS) algorithm, which iteratively improves sub-problem solutions, has been shown to solve large-scale distributed satellite observation scheduling with limited computation and communication zilberstein2025decentralized. We extend the NSS algorithm to the dynamic DCOP setting (DDCOP) to perform scalable and effective dynamic observation scheduling.
Dynamic distributed constraint optimization problems (DDCOP) lass2008dynamic extend DCOPs to capture problem changes. A standard DDCOP is composed of a sequence of $T$ static DCOPs where an optimal solution is obtained by solving each of the $T$ DCOPs optimally. DDCOP solutions are inherently online algorithms as a system reacts to changes. Most work has adapted common DCOP algorithms to dynamic variations that inherit computational complexities mailler2005comparing; faltings2005superstabilizing; khanna2009efficient; billiau2012sbdo; zivan2015distributed; yeoh2015incremental; dynamicdsa.
When it comes to the application of DDCOPs to satellite operations, there are limitations with the standard definition. When solving static DCOPs, it is possible to assume that solutions are found prior to the execution horizon. However, DDCOPs cannot always make this assumption as dynamics may occur during the execution horizon. This is true for satellite operations; utility is obtained by taking an observation, not scheduling one. Therefore, when the problem changes so that a requested task is removed from the problem, having it scheduled prior to execution may not be optimal. In addition, due to consumptive resources, the starting state of subsequent DCOPs is driven by previous solutions.
3. Problem Definition
3.1. Multi-Satellite Constellation Observation Scheduling
We outline the Multi-Satellite Constellation Observation Scheduling Problem (COSP) and discuss the related challenges. COSP is defined by the following sets.
  • $H=[h_{s},h_{e}]$: the scheduling horizon.
  • AA: the set of agents in which each agent is a satellite in the constellation.
  • $\mathcal{T}$: the set of point targets on Earth defined by a latitude and longitude.
  • $R$: the set of requests where each request is defined by the target to observe, $\tau\in\mathcal{T}$, and when in the scheduling horizon to observe, $h\subset H$. Note that we index elements of a request $r$, such as the horizon, with the notation $h(r)$ and use this notation consistently for other variables.
  • $S=\bigcup_{a\in A}S_{a}$: the set of tasks (also referred to as observations) where each $S_{a}$ corresponds to the tasks of agent $a$. A task $s\in S_{a}$ is defined by the request being satisfied, $r\in R$, the interval required to schedule the task, $h\subset h(r)$, and the data volume required to take the observation, $m\in\mathbb{R}^{+}$.
  • $\mathcal{X}=\bigcup_{a\in A}\mathcal{X}_{a}$: the set of Boolean decision variables where each $\mathcal{X}_{a}$ corresponds to the variables of agent $a$. For each $s\in S_{a}$ we define the Boolean decision variable $x\in\mathcal{X}_{a}$ where $x=1$ iff agent $a$ schedules task $s$.
  • $D=\bigcup_{a\in A}D_{a}$: the set of downlinks where each $D_{a}$ corresponds to the downlinks of agent $a$. A downlink is defined by the maximum data volume downlinked, $m\in\mathbb{R}^{+}$, and the time interval of the downlink, $h\subset H$. We assume that all downlinks are mandatory.
  • $C=\bigcup_{a\in A}C_{a}$: the set of constraints for each agent. Each agent is constrained by processing and data volume. An agent cannot execute two tasks at once and tasks cannot overlap with downlinks. An agent must also never exceed its memory capacity and all observations acquired must be downlinked at the earliest opportunity. Formally,
    $C_{a}=C_{D_{a}}\cup C_{S_{a}}.$
    We define
    $C_{D_{a}}=\bigcup_{d\in D_{a}}c_{d}$
    where
    $c_{d}=\sum_{s\in S_{a}^{d}}x(s)\cdot m(s)\leq\textsc{min}(m(a),m(d)).$
    The set $S_{a}^{d}$ contains the possible tasks for which the soonest downlink window in the future is $d$. The value $m(a)$ denotes the memory capacity of agent $a$. We define
    $C_{S_{a}}=\bigcup_{s,s^{\prime}\in S_{a}}c_{s,s^{\prime}}$
    where
    $c_{s,s^{\prime}}=\left[x(s)\cdot x(s^{\prime})+\mathbb{I}(h(s)\cap h(s^{\prime})\neq\emptyset)\leq 1\right].$
The objective of COSP is to maximize the number of requests satisfied subject to the constraints. A request is satisfied if a single observation for that request is completed. An optimal assignment of variables $\mathcal{X}^{*}$ is defined as
$\mathcal{X}^{*}=\arg\max_{\mathcal{X}}\mathcal{F}(\mathcal{X})$
where
$\mathcal{F}(\mathcal{X})=\sum_{r\in R}\left[1-\prod_{x\in\mathcal{X}_{r}}(1-x)\right].$
Here, $\mathcal{X}_{r}$ is the set of variables such that $x=x(s)$ and $r(s)=r$.
COSP has been shown to be a challenging problem for DCOP methods. Typical COSP instances have hundreds of agents and thousands of requests, leading to millions of decision variables. In addition, the constraint graph of COSP has high degrees on the order of $\Omega(|A|\cdot|R|)$. The constraint graph is also assumed to be only locally known to an agent. An agent $a$ is oblivious to all tasks $s\notin S_{a}$ and variables $x\notin\mathcal{X}_{a}$. This is not consistent with standard DCOPs in which agents know all neighboring variables/agents in the constraint graph fioretto2018distributed. In COSP, it is assumed that agents know the existence of all other agents but have no knowledge of the variables of other agents. These factors make existing DCOP approaches that rely on agents communicating with neighboring agents in the constraint graph both computationally challenging due to the high degrees in the graph and inapplicable since we cannot assume agents know which agents they share constraints with.
3.2. Dynamic Multi-Satellite Constellation Observation Scheduling
We now present the Dynamic Multi-Satellite Constellation Observation Scheduling Problem (DCOSP) and discuss how previous challenges from COSP transfer to DCOSP.
A DDCOP consists of a set of $T$ sequential DCOPs. We define DCOSP similarly as a set of COSP instances. Let $\delta_{t}$ be the COSP instance at time $t$. We then define the DCOSP, $\delta$, as $\delta=\{\delta_{t}\}_{t=0}^{T}$. We assume that the agents have no prior knowledge of when or how the problem might change and must act reactively. We refer to the requests and variables of $\delta_{t}$ as $R^{\delta_{t}}$ and $\mathcal{X}^{\delta_{t}}$. Note that $\delta_{i}$ depends on $\delta_{j}$ for $j<i$ since these prior DCOPs will determine the starting state of $\delta_{i}$. For example, resource expenditure affects both current and future solutions. We assume that there is a globally known horizon for a DCOSP instance, $h(\delta)=[h_{s}(\delta),h_{e}(\delta)]$ and that the horizon of each COSP instance is $h(\delta_{t})=[h_{s}(\delta_{t}),h_{e}(\delta)]$ where $h_{s}(\delta_{t})\in h(\delta)$.
The utility of DCOSP is not the same as a typical DDCOP. In a typical DDCOP, the utility is defined as the sum of the utility functions of the individual DCOPs, which means that an optimal solution is obtained by solving each DCOP optimally in sequence. However, this formulation does not adequately capture DCOSP utility. We define the utility of DCOSP to be the number of requests that are satisfied, where satisfaction is determined by an observation for a request being executed. This utility rewards completing a task rather than just scheduling one. Due to the online nature of DCOSP, the scheduling horizon overlaps the execution horizon. Therefore, scheduling an observation does not guarantee that it will be executed. Consider that a task for request $r$ is scheduled at time $t_{0}$ in $\delta_{t_{0}}$ to be executed at time $t_{i}$. If the task is then unscheduled at some $\delta_{t_{j}}$ where $0<j<i$ then $r$ will not be satisfied despite having a task scheduled in COSP instance $\delta_{t_{0}}$.
To formally define this utility, we provide some useful definitions. Let $\bar{h}(\delta_{t})$ be the unknown execution horizon of $\delta_{t}$. We visually show $\bar{h}(\delta_{t})$ in Figure 1. This is the horizon for which the problem is static and is defined by $\delta_{t}$. Formally,
$\bar{h}(\delta_{t})=\begin{cases}[h_{s}(\delta_{t}),h_{s}(\delta_{t+1})]&\text{if}\penalty 10000\ t<T\\ [h_{s}(\delta_{t}),h_{e}(\delta_{t})]&\text{else (i.e. $t=T)$}.\end{cases}$
We then define the proposition $\mathsf{executed}(x)$ as
$\mathsf{executed}(x)=x\cdot\mathbb{I}[\exists t\penalty 10000\ x=x(s)\wedge h(s)\cap\bar{h}(\delta_{t})\neq\emptyset].$
The above is equal to one if and only if $x=1$, meaning task $s$ was scheduled, and the horizon of $s$ occurred during the time when the problem is static as defined by $\delta_{t}$. This means that the task was scheduled when it was executed. Using this proposition, we can define the utility of DCOSP for an assignment of variables $\mathcal{X}^{\delta}$ over all time steps.
$\mathcal{F}(\mathcal{X}^{\delta})=\sum_{r\in R^{\delta}}\left[1-\prod_{x\in\mathcal{X}_{r}^{\delta}}(1-\mathsf{executed}(x))\right]$
where $R^{\delta}=\bigcup_{t=0}^{T}R^{\delta_{t}}$ and $\mathcal{X}_{r}^{\delta}=\bigcup_{t=0}^{T}\mathcal{X}_{r}^{\delta_{t}}$. $R^{\delta}$ is the set of all requests in DCOSP $\delta$ and $\mathcal{X}_{r}^{\delta}$ are all tasks for request $r$ over $\delta$. This utility function rewards observations that are completed, not just scheduled.
The same challenges of solving COSP are multiplied when solving DCOSP. Solving a single instance of COSP is computationally challenging for most DCOP methods. Solutions that are linear in the maximum degree of the constraint graph suffer due to COSP instances having degrees $\Omega(|A|\cdot|R|)$. We again assume that the constraint graph is only partially known to an agent. Therefore, solutions to DCOSP need to conform to this assumption that cross-agent edges in the constraint graph are unknown and to be efficient in computation and communication.
Figure 1. We illustrate $\bar{h}(\delta_{t})$; the line spans the entire horizon of the DCOSP $\delta$ where $\bar{h}(\delta_{t})$ shows the unknown interval the problem is static as defined by $\delta_{t}$.
4. Algorithms
4.1. Obtaining an Optimal Solution to DCOSP
There is no clear online algorithm for computing an optimal DCOSP solution. This is due to agents having no prior knowledge of problem dynamics and optimally solving each individual COSP instance not necessarily constituting an optimal DCOSP solution. However, we can obtain an optimal solution by collapsing DCOSP into a single DCOP. This relies on using omniscient knowledge of the problem dynamics and is therefore not feasible in actuality.
We can collapse a DCOSP $\delta$ into a DCOP $\delta^{\prime}$ to reason only about the observations that matter. The following are the key set constructions of $\delta^{\prime}$:
$S^{\delta^{\prime}}=\{s\in S^{\delta}\penalty 10000\ |\penalty 10000\ \exists_{t}\penalty 10000\ h(s)\cap\bar{h}(\delta_{t})\neq\emptyset\}\text{ and}$
$\mathcal{X}^{\delta^{\prime}}=\{x\in\mathcal{X}^{\delta}\penalty 10000\ |\penalty 10000\ \exists_{t}\penalty 10000\ x=x(s)\wedge h(s)\cap\bar{h}(\delta_{t})\neq\emptyset\}.$
By construction, $\mathsf{executed}(x)=x$ if and only if $x\in\mathcal{X}^{\delta^{\prime}}$ and $\mathsf{executed}(x)=0$ otherwise. Therefore, we obtain the following equivalence.
$\mathcal{F}(\mathcal{X}^{\delta})=\sum_{r\in R^{\delta}}\left[1-\prod_{x\in\mathcal{X}_{r}^{\delta}}(1-\mathsf{executed}(x))\right]$
$\sum_{r\in R^{\delta}}\left[1-\prod_{x\in\mathcal{X}_{r}^{\delta^{\prime}}}(1-x)\right]$
$\mathcal{F}(\mathcal{X}^{\delta^{\prime}}).$
Solving $\delta^{\prime}$ optimally results in an optimal solution to the DCOSP $\delta$. By definition, $\delta^{\prime}$ is a static COSP instance. Therefore, we can employ any complete algorithm to obtain an optimal solution to $\delta$ by solving $\delta^{\prime}$. Although the omniscient solver is not deployable, it serves as a key performance upper bound for benchmarking practical, online algorithms such as D-NSS. However, even solving a single DCOP is NP-Hard adopt. Therefore, for large DCOSP instances, we use an incomplete solver to obtain a near-optimal solution since any exponential time search is infeasible.
4.2. Dynamic Incremental Neighborhood Stochastic Search
In this section, we present the Dynamic Incremental Neighborhood Stochastic Search (D-NSS) algorithm shown in Algorithm 1. Due to the scale of DCOSP and the computational constraints of satellites, we require DDCOP algorithms that are both efficient and can reason about the dynamic nature of the problem. D-NSS extends NSS to address the drawbacks of general DDCOP algorithms. NSS is an iterative algorithm where at each iteration agents stochastically update their variable assignments based on the assignments of agents they communicate with. This iterative procedure is shared with other algorithms such as DSA. However, NSS relies on a decomposition heuristic $\Upsilon:A\times S\rightarrow\{0,1\}$ to generate a subproblem $\mathcal{N}$ for neighborhoods of agents to solve. This subproblem is a smaller DCOP consisting of requests $R_{\mathcal{N}}$ and agents $A_{\mathcal{N}}$. D-NSS continually computes subproblems, repairs local solutions between problem instances, and reasons about prior scheduling and execution in the search and repair phases.
Algorithm 1 Dynamic Incremental Neighborhood Stochastic Search (D-NSS) for agent $a$
Input: $sched,R,A,S_{a},C_{a},\Upsilon,maxIters$
Output: Schedule for agent $a$
[…]
4:return schedsched
We leverage the Geometric Neighborhood Decomposition (GND) heuristic to create these subproblems every time dynamics occur zilberstein2025decentralized. GND efficiently allocates requests to neighborhoods of agents based on the geometry of the constellation and has been shown to effectively partition COSP instances. For completeness, we provide an outline of GND. GND was first introduced with the NSS algorithm and we refer the reader to prior work for a full presentation of the heuristic zilberstein2025decentralized.
GND leverages the orbital geometry of the satellite constellation to hierarchically partition requests to neighborhoods of agents. Let $K$ be the set of orbital planes that define a satellite constellation. For every request, an agent computes the supply from all orbital planes and adds these to estimate the total number of agents with overflights for the request. The supply can also be thought of as an estimate of degrees in the constraint graph. Iterating through requests in ascending order of supply, a request gets assigned to the $n$ neighborhoods with the highest ratio of supply to temporal conflicts. Temporal conflicts are counts of other requests already allocated to a neighborhood that overlap in time with a given request. Finally, within a neighborhood, requests are further subdivided to agents based on biases towards specific tiles on Earth. GND with $n$ degrees of incompleteness is denoted GND($n$).
Algorithm 2 repair for agent $a$
Input: $sched,R,S_{a},C_{a}$
Output: Repaired Schedule for agent $a$
1:for $s\in sched$ do
2:  if $r(s)\notin R$ then
3:   removeFromSchedule($sched,s$)
4:shuffle $S_{a}$
5:for $s\in S_{a}$ do
6:  if $s$ satisfies $C_{a}\penalty 10000\ \wedge\penalty 10000\ r(s)\notin sched$ then
7:   addToSchedule($sched,s$)
8:return schedsched
Algorithm 3 D-NSS-Search for agent $a$
Input: $sched,R_{\mathcal{N}},A_{\mathcal{N}},S_{a},C_{a},maxIters$
Output: Schedule for agent $a$
1:while $i<maxIters\wedge$ not converged do
2:  com_out,com_incom\_out,com\_in $\textsc{message}(A_{\mathcal{N}},$ schedsched)
3:  shuffle $R_{\mathcal{N}}$
4:  for $r\in R_{\mathcal{N}}$ do
5:   assigned $\textsc{stochasticUpdate}(r,sched,com\_in)$
6:   if assigned $\textsc{true}$ then
7:    schedsched $\textsc{schedule}(r,sched,S_{a})$
8:  $i\leftarrow i+1$
9:return schedsched
D-NSS restarts the search phase when changes are initiated in the problem. A key procedure is the repair function that repairs previously computed solutions to leverage assignments of unchanging variables shown in Algorithm 2. Repairing consists of removing all tasks that are no longer in the current problem instance and greedily inserting new tasks in random order. Note that we can also remove from an agent’s schedule all requests that have been previously executed by agents in the same neighborhood. One benefit of the D-NSS algorithm is that it can leverage different repair procedures. We use the random repair function for two main reasons. Random initialization is the standard for DSA and NSS variants in static domains dsa as it promotes diverse solutions and D-NSS is designed to be as lightweight as possible. Computation-intensive repair procedures counteract the efficiency of the algorithm. We show later on that performing random repair improves the quality and efficiency of D-NSS on DCOSP instances.
After each agent repairs its solution, all agents synchronously begin the iteration phase, D-NSS-Search, to fine-tune the repaired solutions. D-NSS-Search extends the search phase of NSS to account for tasks that have just been scheduled versus ones that have been executed. We detail the following sub-procedures.
  • $\textsc{computeSubProblem}(a,A,R,S_{a},\Upsilon)$. This function computes the neighborhood of agent $a$ and the subset of requests for that neighborhood using the decomposition heuristic $\Upsilon$. We use $\Upsilon=$ GND. GND produces neighborhoods that are reflexive and transitive.
  • $\textsc{message}(A_{\mathcal{N}},sched$). This function defines the message exchange between agents in a neighborhood. Each agent $a$ sends to each other agent in $A_{\mathcal{N}}\setminus\{a\}$ the subset of $R_{\mathcal{N}}$ that it has executed a task for already and the subset that it has scheduled in the previous iteration via the variable com_outcom\_out. The resulting data structure com_incom\_in contains the satisfaction information for the neighborhood.
  • $\textsc{stochasticUpdate}(r,sched,com\_in)$. This function computes the assignment of an agent and a request based on the neighborhood’s communication. An assignment refers to if an agent should attempt to schedule a specific request. Let $W$ be the count of agents that scheduled or executed $r$ in the previous iteration, the probability $P_{u}$ be a hyperparamater, $\textsf{executed}(r)$ be the predicate that $r$ was executed already, and $\textsf{assigned}(a,r)$ be the predicate that $a$ is assigned to $r$. An agent computes the probability of assigning to $r$ in the next iteration, P(a,r|com_in)P(a,r|com\_in) using the update scheme from Table 1. For example, according to com_incom\_in, if request $r$ has not been executed, agent $a$ is not assigned to it, and $W\geq 1$, agent $a$ will always remain unassigned to $r$. This update scheme extends the static NSS update schemes to account for dynamic scheduling and execution. If a task for a request has already been executed, then all agents should unassign. Note that $\textsf{executed}(r)\implies W\geq 1$.
    $\mathsf{executed}(r)$ $\neg\mathsf{executed}(r)$ $\neg\mathsf{executed}(r)$
    $\neg\mathsf{assigned}(a,r)$ $\mathsf{assigned}(a,r)$
    $W=0$ N/A 1 $1-P_{u}$
    $W\geq 1$ 0 0 $1/W$
    Table 1. Stochastic assignment update scheme for agent $a$. The table values denote the probability that agent $a$ assigns to request $r$ based on com_incom\_in, P(a,r|com_in)P(a,r|com\_in).
  • $\textsc{schedule}(r,sched,S_{a})$. This function tries to schedule a task for request $r$ given the current schedule. If a task for $r$ satisfies $C_{a}$ it is inserted into the schedule. Otherwise, the scheduler may remove a single task from the schedule to satisfy the constraints. The task with the closest start time to the new task is used as a heuristic for removal. Agent $a$ remains assigned to a removed task and can attempt to re-schedule it in subsequent iterations. Task removal enables the search phase to overcome local maxima. Note that vanilla DCOSP considers all requests of equal priority and there is no temporal flexibility in the start or end times of tasks. However, DCOSP and the scheduling procedure are amenable to these extensions.
Without prior knowledge of the problem dynamics, it is difficult for any online algorithm to reason about which scheduled tasks will be executed. Prioritizing requests that are earlier in the horizon may be more likely not to be removed and yield reward. However, expending resources early on in the horizon results in fewer resources available to handle problem dynamics later. Although D-NSS currently focuses on reactive repair, it can be extended to integrate proactive scheduling through predictive models, which is a subject of future work.
We introduce variables to analyze the complexity of D-NSS. Let $L$ be the maximum size of a satellite’s schedule, $A_{\mathcal{N}}$ be the largest set of agents in a sub-problem, and $R_{\mathcal{N}}$ be the largest set of requests in a sub-problem. In general $L{\ll}|R_{\mathcal{N}}|{\ll}|R|$ since $L$ is constrained by resources and time. Sub-problem computation via GND has a time complexity of $O(|R_{\mathcal{N}}|)$ and uses no communication. The repair procedure is individually computed by an agent and takes $O(|R_{\mathcal{N}}|\log L)$ time and again uses no communication. D-NSS inherits from NSS a computation and communication complexity of $O(|A_{\mathcal{N}}|{\cdot}|R_{\mathcal{N}}|)$ during an iteration.
4.3. Baseline Algorithms
We evaluate several baseline algorithms in addition to D-NSS. The naive alternative to D-NSS is to recompute a solution from scratch every time problem dynamics occur. We refer to this procedure as 0-NSS. 0-NSS has the same theoretical complexities as D-NSS. We also evaluate dynamic variations of the Distributed Stochastic Search (DSA) algorithm dsa; shreya. D-DSA uses Algorithm 2 to repair DSA solutions. Likewise, 0-DSA runs DSA from scratch every time the problem changes. These DSA variants have a complexity of $O(|A|{\cdot}|R|)$ per iteration. Dynamic DSA is one of the most lightweight DDCOP solvers and can be deployed without violating the constraint graph assumptions of DCOSP. Other DDCOP algorithms such as a dynamic variation of MGM mgm, would also incur a computation and communication complexity of $O(|A|{\cdot}|R|)$ per iteration. Due to the scale of the problem instances and the constraints of satellite computation and communication, we are unable to evaluate any exponential-time algorithms.
We include two baseline non-communication-reliant algorithms: random and greedy. Both of these algorithms construct a schedule for a satellite without reasoning about other agents. These algorithms transition schedules after the dynamics occur by removing redundant tasks. The random solver randomly inserts tasks into an agent’s schedule while the greedy algorithm orders tasks by their start time and iteratively inserts them into a schedule in a single pass. Although simple, a greedy solver is comparable to deployed planners on spacecraft m2020-planner-astra2022.
Finally, we use the construction outlined previously to obtain optimal and near-optimal solutions. For small problems, we obtain an optimal solution by constructing a static COSP instance and solving it with a centralized branch and bound. For large problems, computing an optimal solution is not computationally feasible, so we lower bound the optimal solution using Squeaky Wheel Optimization (SWO) swo, an incomplete centralized solver.
5. Experiments
5.1. Setup
In this section, we outline the experimental setup including constellations and dynamic observation campaigns.
5.1.1. Satellite Constellations
We evaluate two constellations modeled after operational low Earth orbit constellations planet. The Planet constellation is modeled on the Dove constellation from Planet Labs. This constellation is composed of two near sun-synchronous orbital planes at $95^{\circ}$ inclinations each composed of $95$ satellites with an additional two orbital planes at $52^{\circ}$ inclinations each with $5$ satellites. The Walker constellation is motivated by the Skysat constellation from Planet Labs. This constellation has $6$ orbital planes with $14$ satellites each at an $88^{\circ}$ inclination and an overlay of $2$ orbital planes at a $51.6^{\circ}$ inclination with $12$ satellites. Each satellite has a memory capacity of $125$ GB and a single sensor that can slew to $60^{\circ}$ and $45^{\circ}$ off-nadir for the Planet and Walker constellations respectively. Figure 2 depicts these two constellations.
Figure 2. The satellite constellations: Planet (left) and Walker (right). Dots show satellites in an orbital plane.
5.1.2. Downlinks
A $62.5$ MB/s constant bit stream models the satellite downlink during visibility periods with a ground station. We incorporate two ground stations: the ASF Near Space Network Satellite Tracking Ground Station and the Guam Remote Ground Terminal System.
5.1.3. Dynamic Observation Request Campaigns
The target set, $\mathcal{T}$, is $634$ globally distributed cities. Dynamic campaigns consist of periodic requests of these targets. For large problem instances, a periodicity is uniformly sampled from the range $[5,12]$. A target with periodicity $p$ is requested to be observed once within $p$ evenly spaced intervals. For small problem instances, we fix the periodicity at $3$. The scheduling horizon is set to be one day. However, the start of the horizon is randomly initialized. An observation’s memory consumption is sampled from a normal distribution with mean $50$ MB and standard deviation $10$ MB. The time interval required to execute an observation is $63$ seconds which accounts for imaging, slewing, and processing.
To generate dynamics, we select a volatility parameter $v$. This determines the number of changes during the horizon. The time of the dynamics is uniformly distributed over the final $1-\frac{2}{3\cdot v}$ of the scheduling horizon. This ensures that dynamics do not occur directly after the start of the horizon. The set of starting requests is randomly selected from the request set and initialized to one-third of the total size. We select from the remaining requests a random set of $\frac{2}{3\cdot v}$ to add. From the active requests, we randomly remove $\frac{1}{3\cdot v}$ of them. We enforce that once a request is removed, it is not added back, and requests are not changed after their execution horizon starts. We sample $v\in[3,5]$ uniformly.
5.1.4. Hyperparameters and Execution Environment
The stochastic update of NSS and DSA relies on the probability, $P_{u}$ that determines when an agent should unassign from a task. We fix this value at $0.7$, as published in previous work shreya; zilberstein2025decentralized. The $maxIters$ is set to $20$. For the GND heuristic, we set all hyperparameters as described in previous work, and specifically use the GND(2) heuristic zilberstein2025decentralized. We use a random seed of $2005$ for the initial scenario generation. When evaluating $N$ scenarios, we increment this seed for generation of each subsequent simulation. The random seed used in the repair procedure is initialized to $1$. The NSS and DSA algorithms are seeded with a value of $1234$, and the random seed for GND is set to $2$. The random scheduling algorithm uses a seed of $2023$. All experiments are executed using Java 19 on a MacBook Pro 16 laptop with an M2 Max processor (12-core CPU and 38-core GPU) and 64 GB of RAM.
5.2. Results on Small Problem Instances
(a) Planet constellation satisfaction %
(b) Planet constellation message volume
(c) Planet constellation execution time
(d) Walker constellation satisfaction %
(e) Walker constellation message volume
(f) Walker constellation execution time
(g) Satisfaction % of iterative algorithms for the Planet constellation.
(h) Satisfaction % of iterative algorithms for the Walker constellation.
Figure 3. Results of 10 large-scale simulations for the Planet (top row) and Walker (middle row) constellation. We report the average satisfaction, total message volume, and per-agent runtime. Note the log scale for message volume and runtime. We also show the solution quality across a fixed number of iterations for iterative algorithms for instances with $v=5$ (bottom).
Algorithm Opt. Gap (%) Time (ms) Messages (KB)
Random 2.530 <<1 0
Greedy 8.373 <<1 0
D-NSS 1.867 <<1 7.3
0-NSS 2.590 <<1 13.2
D-DSA 4.217 5.3 980.6
0-DSA 3.795 4.7 718.4
Table 2. Results on 10 dynamic problems for the Planet constellation (up to 500 requests).
Algorithm Opt. Gap (%) Time (ms) Messages (KB)
Random 14.945 <<1 0
Greedy 15.604 <<1 0
D-NSS 0.142 5.2 240.2
0-NSS 0.480 6.1 400.0
D-DSA 1.165 58.2 10,459.5
0-DSA 1.215 54.0 7,789.0
Table 3. Results on 10 dynamic problems for the Walker constellation (up to 1000 requests).
We are able to obtain an optimal solution for small problem instances using the omniscient offline algorithm. We solve this DCOP with a centralized branch and bound to obtain an optimal schedule for each satellite. For the Planet constellation, we solve campaigns of up to $500$ requests. For the Walker constellation, this increases to $1000$ requests. In addition to having fewer agents, the Walker constellation geometry under-constrains small problems, making finding optimal solutions faster. An optimal solver does not consistently terminate for larger problems. We report the average gap in satisfaction percentage to the optimal solution for $10$ dynamic small problem instances in Tables 2 and 3.
These results support the theoretical analysis of D-NSS and demonstrate near-optimal performance. D-NSS outperforms all baselines, including D-DSA, 0-DSA, and 0-NSS, and does so while using less computation and communication. Notably, compared to DSA variants, D-NSS finds better solutions using an order of magnitude less computation and up to two orders of magnitude less communication. This is due to both the lower theoretical complexity per iteration and the faster convergence of D-NSS.
5.3. Results on Large Problem Instances
We evaluate the algorithms on realistic, large-scale, dynamic scenarios with thousands of requests. Figure 3 shows the results for the Planet and Walker constellations. We report the total utility of each algorithm in Figures 3(a) and 3(d), the total message volume in Figures 3(b) and 3(e), and the average per-agent execution time in Figures 3(c) and 3(f). Note the log scale in the figures for message volume and runtime.
In terms of solution quality, D-NSS and D-DSA achieve close to the optimal lower-bound solution. D-NSS outperforms both 0-NSS and 0-DSA as well as the greedy and random baselines. Crucially, D-NSS achieves high solution quality while having a lower total message volume and execution time compared to D-DSA, 0-DSA, and 0-NSS. Both D-NSS and 0-NSS use an order of magnitude fewer messages and computation times than their DSA counterparts. Since the optimal lower bound algorithm is computed centrally, the message volume corresponds to the schedules a ground station would have to uplink.
We also evaluate the stability and convergence of D-NSS compared to D-DSA, 0-DSA, and 0-NSS. Figures 3(g) and 3(h) show the average solution quality over iterations of the algorithms during large-problem instances with $v=5$. We fix the number of iterations for all algorithms to show the relative performance. Clearly, D-NSS and D-DSA are much more stable, reacting much less volatility to problem dynamics, illustrated by the sharp drops in solution quality of 0-NSS and 0-DSA. Notably, D-NSS is very stable; even in the iteration directly after problem dynamics, D-NSS repairs solutions effectively to maintain or improve solution quality. This supports computing and repairing sub-problems rather than global solutions like D-DSA. In addition, the stability of D-NSS leads to a quick convergence in practice, which drives the efficiency of the algorithm.
6. Conclusion
Large problems remain a challenge for DDCOP algorithms due to their computational and communication complexities. Even linear time and messaging solutions may not be feasible when constraint graphs are highly connected or agents have limited compute. These challenges apply to DCOSP. We present the D-NSS algorithm, a decomposition-based algorithm that is efficient in time and message complexity and can scale to problems much larger than previous approaches. We show that D-NSS stabilizes quickly and uses an order of magnitude fewer messages and compute times compared to DDCOP baselines. Despite no quality guarantees, we show that D-NSS computes near-optimal solutions.
D-NSS generalizes to many DDCOP problems. While we leverage the GND heuristic, the D-NSS framework can accommodate alternative heuristics, including learned or adaptive ones, enabling broader applicability across DDCOP domains. Some examples include multi-agent path finding mapf, mobile sensor teams ca-dcop-mst, and UAV coordination uavs-coord.
In future work, we can extend D-NSS to reason about future dynamics. We can construct or learn heuristics that prioritize tasks that are more likely to yield utility when executed. Alternatively, we can extend DCOSP to be a Proactive DCOP (PDDCOP) pddcop where we have priors on problem dynamics, and agents can generate robust offline solutions. However, PDDCOPs require a model of problem dynamics which may not be obtainable in practice.
DCOSP solutions will be demonstrated in operational scenarios beginning in 2026. One example is NASA’s FAME demonstration fame-spaceops-2025 which centers around a federated observation system zilberstein-spaceops-2025. The FAME demonstration consists of more than 60 participating Earth-observing spacecraft that will coordinate their measurements to optimize observation completion across dynamic scenarios. Through onboard control, satellites will be able to react faster and with more precision to time-sensitive events such as natural disasters. These observations will provide novel measurements of scientific processes and crucial, up-to-date information for human operators. This work has laid the foundation for dynamic, large-scale, distributed onboard scheduling of Earth-observing satellites.
{acks}
Portions of the research were carried out at the Jet Propulsion Laboratory, California Institute of Technology, under a contract with the National Aeronautics and Space Administration (80NM0018D0004). Government sponsorship acknowledged. This material is based upon work supported by the National Science Foundation Graduate Research Fellowship Program under Grant No DGE2140739. Any opinions, findings, and conclusions or recommendations expressed in this material are those of the author(s) and do not necessarily reflect the views of the National Science Foundation.